CLNS 94/1304 CLEO 94-22

Observation of excited charmed baryon states decaying to Λ_{c}^{+}π^{+}π^{−}

K.W. Edwards,^{1} M. Ogg,^{1} A. Bellerive,^{2} D.I. Britton,^{2} E.R.F. Hyatt,^{2} D.B. MacFarlane,^{2} P.M. Patel,^{2} B. Spaan,^{2} A.J. Sadoff,^{3} R. Ammar,^{4} P. Baringer,^{4} A. Bean,^{4} D. Besson,^{4} D. Coppage,^{4} N. Copty,^{4} R. Davis,^{4} N. Hancock,^{4} M. Kelly,^{4} S. Kotov,^{4} I. Kravchenko,^{4} N. Kwak,^{4} H. Lam,^{4} Y. Kubota,^{5} M. Lattery,^{5} M. Momayezi,^{5} J.K. Nelson,^{5} S. Patton,^{5}

R. Poling,^{5} V. Savinov,^{5} S. Schrenk,^{5} R. Wang,^{5} M.S. Alam,^{6} I.J. Kim,^{6} Z. Ling,^{6} A.H. Mahmood,^{6} J.J. O’Neill,^{6} H. Severini,^{6} C.R. Sun,^{6} F. Wappler,^{6} G. Crawford,^{7} C. M. Daubenmier,^{7} R. Fulton,^{7} D. Fujino,^{7} K.K. Gan,^{7} K. Honscheid,^{7} H. Kagan,^{7} R. Kass,^{7} J. Lee,^{7} M. Sung,^{7} C. White,^{7} A. Wolf,^{7} M.M. Zoeller,^{7} F. Butler,^{8} X. Fu,^{8} B. Nemati,^{8} W.R. Ross,^{8} P. Skubic,^{8} M. Wood,^{8} M. Bishai,^{9} J. Fast,^{9} E. Gerndt,^{9} J.W. Hinson,^{9} R.L. McIlwain,^{9} T. Miao,^{9} D.H. Miller,^{9} M. Modesitt,^{9} D. Payne,^{9} E.I. Shibata,^{9} I.P.J. Shipsey,^{9} P.N. Wang,^{9} M. Battle,^{10} J. Ernst,^{10} L. Gibbons,^{10}

Y. Kwon,^{10} S. Roberts,^{10} E.H. Thorndike,^{10} C.H. Wang,^{10} J. Dominick,^{11} M. Lambrecht,^{11} S. Sanghera,^{11} V. Shelkov,^{11} T. Skwarnicki,^{11} R. Stroynowski,^{11} I. Volobouev,^{11} G. Wei,^{11}

P. Zadorozhny,^{11} M. Artuso,^{12} M. Gao,^{12} M. Goldberg,^{12} D. He,^{12} N. Horwitz,^{12} G.C. Moneti,^{12} R. Mountain,^{12} F. Muheim,^{12} Y. Mukhin,^{12} S. Playfer,^{12} Y. Rozen,^{12} S. Stone,^{12} X. Xing,^{12} G. Zhu,^{12} J. Bartelt,^{13} S.E. Csorna,^{13} Z. Egyed,^{13} V. Jain,^{13} D. Gibaut,^{14} K. Kinoshita,^{14} P. Pomianowski,^{14} B. Barish,^{15} M. Chadha,^{15} S. Chan,^{15} D.F. Cowen,^{15} G. Eigen,^{15} J.S. Miller,^{15} C. O’Grady,^{15} J. Urheim,^{15} A.J. Weinstein,^{15} M. Athanas,^{16} W. Brower,^{16} G. Masek,^{16} H.P. Paar,^{16} J. Gronberg,^{17} C.M. Korte,^{17}

R. Kutschke,^{17} S. Menary,^{17} R.J. Morrison,^{17} S. Nakanishi,^{17} H.N. Nelson,^{17} T.K. Nelson,^{17} C. Qiao,^{17} J.D. Richman,^{17} A. Ryd,^{17} D. Sperka,^{17} H. Tajima,^{17} M.S. Witherell,^{17} M. Procario,^{18} R. Balest,^{19} K. Cho,^{19} W.T. Ford,^{19} D.R. Johnson,^{19} K. Lingel,^{19}

M. Lohner,^{19} P. Rankin,^{19} J.G. Smith,^{19} J.P. Alexander,^{20} C. Bebek,^{20} K. Berkelman,^{20} K. Bloom,^{20} T.E. Browder,^{20}^{∗} D.G. Cassel,^{20} H.A. Cho,^{20} D.M. Coffman,^{20} D.S. Crowcroft,^{20} P.S. Drell,^{20} D.J. Dumas,^{20} R. Ehrlich,^{20} P. Gaidarev,^{20}

M. Garcia-Sciveres,^{20} B. Geiser,^{20} B. Gittelman,^{20} S.W. Gray,^{20} D.L. Hartill,^{20} B.K. Heltsley,^{20} S. Henderson,^{20} C.D. Jones,^{20} S.L. Jones,^{20} J. Kandaswamy,^{20}

N. Katayama,^{20} P.C. Kim,^{20} D.L. Kreinick,^{20} G.S. Ludwig,^{20} J. Masui,^{20} J. Mevissen,^{20} N.B. Mistry,^{20} C.R. Ng,^{20} E. Nordberg,^{20} J.R. Patterson,^{20} D. Peterson,^{20} D. Riley,^{20} S. Salman,^{20} M. Sapper,^{20} F. Würthwein,^{20} P. Avery,^{21} A. Freyberger,^{21} J. Rodriguez,^{21} S. Yang,^{21} J. Yelton,^{21} D. Cinabro,^{22} T. Liu,^{22} M. Saulnier,^{22} R. Wilson,^{22} H. Yamamoto,^{22}

T. Bergfeld,^{23} B.I. Eisenstein,^{23} G. Gollin,^{23} B. Ong,^{23} M. Palmer,^{23} M. Selen,^{23} and J. J. Thaler^{23}

(CLEO Collaboration)

1Carleton University, Ottawa, Ontario K1S 5B6 and the Institute of Particle Physics, Canada 2McGill University, Montréal, Québec H3A 2T8 and the Institute of Particle Physics, Canada

3Ithaca College, Ithaca, New York 14850 4University of Kansas, Lawrence, Kansas 66045 5University of Minnesota, Minneapolis, Minnesota 55455 6State University of New York at Albany, Albany, New York 12222 7Ohio State University, Columbus, Ohio, 43210 8University of Oklahoma, Norman, Oklahoma 73019 9Purdue University, West Lafayette, Indiana 47907 10University of Rochester, Rochester, New York 14627 11Southern Methodist University, Dallas, Texas 75275 12Syracuse University, Syracuse, New York 13244 13Vanderbilt University, Nashville, Tennessee 37235

14Virginia Polytechnic Institute and State University, Blacksburg, Virginia, 24061 15California Institute of Technology, Pasadena, California 91125 16University of California, San Diego, La Jolla, California 92093 17University of California, Santa Barbara, California 93106 18Carnegie-Mellon University, Pittsburgh, Pennsylvania 15213 19University of Colorado, Boulder, Colorado 80309-0390 20Cornell University, Ithaca, New York 14853 21University of Florida, Gainesville, Florida 32611 22Harvard University, Cambridge, Massachusetts 02138 23University of Illinois, Champaign-Urbana, Illinois, 61801

(January 12, 1995)

Abstract

Using data collected by the CLEO II detector, we have observed two states decaying to Λ_{c}^{+}π^{+}π^{−}. Relative to the Λ_{c}^{+}, their mass splittings are measured to be +307.5 ± 0.4 ± 1.0, and +342.2 ± 0.2 ± 0.5 MeV/c^{2}, respectively; this represents the first measurement of the less massive state. These two states are consistent with being orbitally excited, isospin zero Λ_{c}^{+} states.

13.30.Eg, 14.20.Lq

Typeset using REVTEX

∗Permanent address: University of Hawaii at Manoa

Soon after the discovery of charm, several models predicted a rich charmed baryon spectroscopy [1] [2] [3] [4] [5]. As expected in Heavy Quark Effective Theory (HQET) [6], the mass splittings observed in charmed systems, compared with those observed in strange baryons and mesons, were qualitatively consistent with the expected m_{Q}^{−1} (with m_{Q} designating the mass of the heavy quark) dependence of the color-magnetic dipole moment of quarks. This m_{Q}^{−1} dependence of the mass splittings is, in fact, one of the starting points of HQET. In HQET, the main features of hadron spectroscopy may be understood in terms of light quarks moving in the heavy quark’s confinement potential, perturbed by various short-range interactions, as anticipated from one gluon exchange.

In this Letter, we report the observation of two excited charmed baryons, which we will refer to as the Λ_{c}^{∗+}(2625) and the Λ_{c}^{∗+}(2593). The former of these has previously been reported by both the ARGUS [7] and E687 [8] collaborations; a preliminary version of the analysis contained herein, which reported evidence for the lower state, has been previously presented [9]. Evidence that these two excited charmed baryons correspond to the lowest lying spatial excitations of the Λ_{c}^{+} baryon is also given.

The CLEO II detector is discussed in detail elsewhere [10]. This analysis involves mainly the central tracking system, consisting of two precision vertex chambers and a cylindrical wire drift chamber, all inside a 1.5 T axial magnetic field. Outside the drift chamber is a time-of-flight system, which is used, in conjunction with specific ionization measurements in the drift chamber, for particle identification. The data set used for this analysis corresponds to an integrated luminosity of approximately 3 fb^{−1} of e^{+}e^{−} collisions, taken at center of mass energies between 10.52 and 10.58 GeV.

We obtain a Λ_{c}^{+} sample using six decay modes, pK^{−}π^{+}, p\overline{K}^{0}, Λπ^{+}, Λπ^{+}π^{0}, Λπ^{+}π^{−}π^{+} and Σ^{+}π^{+}π^{−} (charge conjugate modes are implicit). For the pK^{−}π^{+} and p\overline{K}^{0} modes we use a combination of the measurements of specific ionization in the drift chamber and timeof-flight information to identify the p and K^{−} candidates. The \overline{K}^{0} mesons were identified through their decay \overline{K}^{0} → K_{s}^{0}, K_{s}^{0} → π^{+}π^{−}, and the Λ candidates by their decay Λ → pπ^{−}. In both cases a secondary vertex is reconstructed from a pair of oppositely charged tracks intersecting at a point well separated from the primary vertex. The pair is identified as a K_{s}^{0}(Λ) candidate if the invariant mass of the π^{+}π^{−}(pπ^{−}) pair is consistent with the K_{s}^{0}(Λ) mass, and if the momentum vector of the candidate K_{s}^{0}(Λ) points back to the primary event vertex. Σ^{+} candidates were found by forming pπ^{0} combinations which are consistent with coming from the decay of a Σ^{+} with a decay point displaced from the primary vertex [11]. Our sample of π^{0} candidates is defined by photon pairs having a measured invariant mass within ±2.5σ of the known π^{0} mass (σ_{m_{π^{0}}} ∼5 MeV/c^{2}); any cluster in the calorimeter with greater than 50 MeV of deposited energy and not matched to the position of charged tracks extrapolated outwards from the central tracking chamber is considered a photon candidate.

We require Λ_{c}^{+} candidates to have an invariant mass within ±20 MeV/c^{2} of the known Λ_{c}^{+} mass. For purposes of background studies, we define two sideband samples for each mode, also ±20 MeV/c^{2} wide, and centered at ±45 MeV/c^{2} from the central Λ_{c}^{+} mass value. Each Λ_{c}^{+} candidate is then combined with all remaining pairs of oppositely charged particles, assumed to be pions, and the mass difference M(Λ_{c}^{+}π^{+}π^{−}) - M(Λ_{c}^{+}) is then calculated. Loose particle identification requirements are imposed on the transition pions in order to suppress any possible backgrounds from D^{∗} → Dγ or D^{∗} → Dπ^{0}, which may produce a low-mass electron-positron pair in the case where a transition photon pair converts, or a transition π^{0}

FIG. 1. Fit to mass difference: M(Λ_{c}^{+}π^{+}π^{−})-M(Λ_{c}^{+}).

undergoes a Dalitz decay. To improve signal to noise, the Λ_{c}^{+}π^{+}π^{−} combinations must satisfy: x_{p} > 0.7, where x_{p} is the ratio of the candidate momentum to the maximum kinematically allowed momentum for e^{+}e^{−} production of charmed particle pairs (x_{p} = p/\sqrt{E_{beam}^{2} − M_{Λ_{c}^{∗+}}^{2}}). The mass difference spectrum for combinations satisfying the above cuts is shown by the data points in Figure 1.

A large peak, visible at a mass difference of approximately 342 MeV/c^{2}, is due to the previously reported decay Λ_{c}^{∗+}(2625) → Λ_{c}^{+}π^{+}π^{−}. Also present is a smaller, but still significant, enhancement at a mass difference of ≈308 MeV/c^{2}. The sum of a polynomial background shape plus two signal functions is used to fit the mass difference distribution. Each signal is described as a spin-0 Breit-Wigner line shape, with a floating intrinsic width, convoluted with a Gaussian of fixed width to parameterize the detector resolution. The width of this Gaussian is taken from Monte Carlo studies. The fitted signal sizes of the upper and lower peaks are 244.6 ± 19.0 and 112.5 ± 16.5 events, respectively. The results of the fits, giving the observed yield, extracted widths, and mass splittings relative to the Λ_{c}^{+}, are given in Table I. The systematic errors on the quoted widths are obtained by observing the variation in width when the event selection criteria are varied, or when the signal plus background parameterizations are varied. The systematic error on the mass difference reflects uncertainties in the overall momentum scale, signal parameterization, and energy loss correction in the detector at low momentum.

In order to verify that the signals are not an artifact of the selection criteria, two checks were performed, using combinations from the Λ_{c}^{+} sidebands, as well as same-sign (Λ_{c}^{+}π^{±}π^{±}) combinations. Neither the mass difference plot for combinations from the Λ_{c}^{+} mass sidebands nor the plot for same-sign combinations shows any departure from smooth behavior, and both are consistent in shape with our background parameterization.

If these states are isospin zero objects, then a strong decay into Λ_{c}^{+}π is prohibited. An electromagnetic decay into Λ_{c}^{+}γ is allowed, and may be competitive with dipion transitions if the intrinsic widths of the states are sufficiently narrow. If, on the other hand, these objects are excited Σ_{c} baryons, in which the ud diquark is in an isospin 1 state, then their dominant

TABLE I. Parameters of fit to Λ_{c}^{∗+} signals. Also shown are results from the ARGUS and E687 experiments.

Decay  Events  Γ_{Λ_{c}^{∗+}}  M(Λ_{c}^{∗+}) − M(Λ_{c}^{+})
(MeV/c^{2})  (MeV/c^{2})
Λ_{c}^{∗+}(2625) → Λ_{c}^{+}π^{+}π^{−}  244.6±19.0  <1.9  342.2±0.2±0.5
Λ_{c}^{∗+}(2593) → Λ_{c}^{+}π^{+}π^{−}  112.5±16.5  3.9_{−1.2−1.0}^{+1.4+2.0}  307.5 ± 0.4 ± 1.0
Λ_{c}^{∗+}(2625) → Λ_{c}^{+}π^{0}  62±50
Λ_{c}^{∗+}(2625) → Λ_{c}^{+}γ  -5±24
Λ_{c}^{∗+}(2593) → Λ_{c}^{+}π^{0}  -44±47
Λ_{c}^{∗+}(2593) → Λ_{c}^{+}γ  11±26

(Previous results)

Λ_{c}^{∗+}(2625) → Λ_{c}^{+}π^{+}π^{−} [7]  45.6±10.1  <3.2  341.5±0.6±1.6 [12]
Λ_{c}^{∗+}(2625) → Λ_{c}^{+}π^{+}π^{−} [8]  39.7±8.7  340.4±0.6±0.3

decay should be to Λ_{c}^{+}π. We have therefore also searched for these resonances in the final states Λ_{c}^{+}π^{0} (neutral pions as defined previously) and Λ_{c}^{+}γ (photons as defined previously, but with the added requirement that the measured energy exceed 200 MeV). Results of our searches in these modes are given in Table I; no signals are found and we therefore quote only upper limits. We determine: \frac{B(Λ_{c}^{∗+}(2593)→Λ_{c}^{+}π^{0})}{B(Λ_{c}^{∗+}(2593)→Λ_{c}^{+}π^{+}π^{−})} < 3.53 and \frac{B(Λ_{c}^{∗+}(2593)→Λ_{c}^{+}γ)}{B(Λ_{c}^{∗+}(2593)→Λ_{c}^{+}π^{+}π^{−})} < 0.98. The corresponding upper limits for the Λ_{c}^{∗+}(2625) are 0.91 and 0.52, respectively. For a charmed baryon just above Λ_{c}^{+}π^{+}π^{−} threshold, the phase space for Λ_{c}^{+}π decay is much more favorable than for Λ_{c}^{+}π^{+}π^{−} decay. The combination of our non-observation of Λ_{c}^{∗+} in the Λ_{c}^{+}π channel, coupled with observation in the Λ_{c}^{+}π^{+}π^{−} channel, supports the interpretation of these states as Λ_{c}^{∗+} baryons rather than Σ_{c}^{∗+} baryons [14].

The presence of substructure in the Λ_{c}^{+}π^{+}π^{−} final state (Σ_{c}π, e.g.) can also provide information on the identities of these states. Of interest is the fraction, f_{Σ_{c}}, of the Λ_{c}^{∗+}(2625) → Λ_{c}^{+}π^{+}π^{−} decay rate which proceeds through an intermediate Σ_{c}. Previous analyses of this final state by the ARGUS [7] and E687 [8] collaborations determined f_{Σ_{c}} to be 0.46±0.14, and <36%, respectively. Figure 2 shows the Λ_{c}^{+}π^{−} - Λ_{c}^{+} mass difference plotted against the Λ_{c}^{+}π^{+} - Λ_{c}^{+} mass difference for our Λ_{c}^{+}π^{+}π^{−} final sample. A cluster of events, corresponding to transitions of the Λ_{c}^{∗+}(2593) decaying to Λ_{c}^{+} through an intermediate Σ_{c}, is prominent at each end of the diagonal band near the lower left corner of the plot. Such clustering is not evident for the band corresponding to the Λ_{c}^{∗+}(2625). To determine the fraction of times that the decay Λ_{c}^{∗+} → Λ_{c}^{+}π^{+}π^{−} occurs through the intermediate state Λ_{c}^{∗+} → Σ_{c}π, we measure the Λ_{c}^{∗+} yield as a function of Λ_{c}^{+}π sub-mass. By subtracting the Λ_{c}^{∗+} yield when the Λ_{c}^{+}π sub-mass is in the Σ_{c} sidebands from the Λ_{c}^{∗+} yield when the Λ_{c}^{+}π mass is consistent with the Σ_{c} mass, we extract the values f_{Σ_{c}^{++}} and f_{Σ_{c}^{0}}, corresponding to the Σ_{c}^{++}π^{−} and Σ_{c}^{0}π^{+} content, respectively, of the final state. Our measurements of these parameters are given in Table II [15].

Also of interest is the fragmentation function of these states compared to that of the ground state Λ_{c}^{+}(2285). The orbitally excited charmed mesons D_{1}(2420) and D_{2}(2460), as well as the orbitally excited D_{sJ} states, have measured fragmentation functions which are noticeably harder than either those of the ground state D and D_{s} pseudoscalar mesons or those of the corresponding vector states [16]. Figure 3 shows the efficiency corrected

FIG. 2. Scatter plot of the mass difference M(Λ_{c}^{+}π^{−}) - M(Λ_{c}^{+}) vs. the mass difference M(Λ_{c}^{+}π^{+})

- M(Λ_{c}^{+}), for events with M(Λ_{c}^{+}π^{+}π^{−})<2650 MeV/c^{2}. Visible in the figure are (1) The vertical band at Λ_{c}^{+}π^{+} mass difference of 168 MeV/c^{2}, corresponding to Σ_{c}^{++} → Λ_{c}^{+}π^{+}; (2) The horizontal band at Λ_{c}^{+}π^{−} mass difference of 168 MeV/c^{2}, corresponding to Σ_{c}^{0} → Λ_{c}^{+}π^{−}; (3) The diagonal band extending from (0.14,0.20) to (0.20,0.14), corresponding to combinations taken from the Λ_{c}^{∗+}(2625) signal region in Figure 1 (squares); (4) The diagonal band extending from (0.14,0.168) to (0.168,0.14), corresponding to combinations taken from the Λ_{c}^{∗+}(2593) signal region in Figure 1 (triangles), within which there are two lobes at the kinematic limits: a) The enhancement in the Σ_{c}^{++} band, just above the threshold for Λ_{c}^{+}π^{−}, and b) The enhancement in the Σ_{c}^{0} band, just above the threshold for Λ_{c}^{+}π^{+}.

FIG. 3. Normalized dN/dx_{p} spectra of the Λ_{c}^{∗+} samples. Shown are the data for the Λ_{c}^{∗+}(2625) (diamonds) and the Λ_{c}^{∗+}(2593) (inverted triangles), with fits to the Peterson function overlaid as the dashed and solid curves, respectively.

TABLE II. Λ_{c}^{∗+} e^{+}e^{−} production and decay results from this analysis and ARGUS[7].

Signal  ϵ_{P}  f_{Σ_{c}^{++}}  f_{Σ_{c}^{0}} (%)  \frac{σ(Λ_{c}^{∗+}→Λ_{c}^{+}π^{+}π^{−})}{σ(Λ_{c}^{+})} (%)
Λ_{c}^{+}(2285) [19]  0.27±0.05
Λ_{c}^{∗+}(2593) 0.057 ± 0.023 ± 0.016 0.36 ± 0.09 ± 0.09 0.42 ± 0.09 ± 0.09 1.44±0.24±0.30
Λ_{c}^{∗+}(2625) 0.065±0.016±0.013 < 0.08 (90% c.l.) < 0.07 (90% c.l.) 3.51±0.34±0.28
Λ_{c}^{∗+}(2625) [7]  0.044±0.018  f_{Σ_{c}^{++}} + f_{Σ_{c}^{0}}=0.46±0.14  4.1±1.0±0.8

fragmentation functions for the Λ_{c}^{∗+} sample. We have fit our data using the fragmentation function parameterization of Peterson [17] in terms of one parameter ϵ_{P}; as the average scaled momentum (x_{p}) approaches 1, ϵ_{P} approaches zero. The values of Peterson ϵ_{P} we extract for the Λ_{c}^{∗+} samples are given in Table II. For purposes of comparison, ϵ_{P} of the ground state Λ_{c}^{+} is also included in the Table. The Λ_{c}^{∗+} baryons clearly display harder momentum spectra. The systematic errors given in our determination of ϵ_{P} are dominated by uncertainties in the absolute tracking efficiency of the soft transition pions.

From the efficiency corrected cross-section, σ(e^{+}e^{−} → Λ_{c}^{∗+}X) · B(Λ_{c}^{∗+} → Λ_{c}^{+}π^{+}π^{−}), we can determine the fraction of Λ_{c}^{+} baryons observed in CLEO which are actually produced as Λ_{c}^{∗+} daughters through Λ_{c}^{∗+} → Λ_{c}^{+}π^{+}π^{−} (using the Peterson function to extrapolate to zero momentum) [18]. We determine these fraction to be 1.44 ± 0.24 ± 0.30% and 3.51 ± 0.34 ± 0.28%, for the Λ_{c}^{∗+}(2593) and Λ_{c}^{∗+}(2625), respectively, as shown in Table II. This is comparable to the fraction of L=0 charmed mesons which are determined to be decay products of orbitally excited D_{1}(2420) or D_{2}^{∗}(2460) mesons [16]. The systematic error in this fraction is dominated by uncertainties in our track reconstruction efficiency.

In the strange baryon sector, the two lowest lying spatial excitations have total spin 1/2 and 3/2 (the Λ(1405) and the Λ(1520), respectively). These belong to a multiplet with the (ud) diquark in a state of S=0, L=0, orbiting the strange quark with relative orbital angular momentum L=1. This orbital angular momentum (which corresponds to the total angular momentum of the light degrees of freedom) combines with the spin of the heavier strange quark to produce two states, one having total spin 1/2 and the other having total spin 3/2. It has been predicted [1–5] that there should be two analogous Λ_{c}^{∗+} states in the charm sector. These orbitally excited Λ_{c}^{∗+} baryons would have preferred decays to Σ_{c}π and Σ_{c}^{∗}π, respectively, much as the orbitally excited strange baryons decay dominantly to Σπ rather than to Λπ^{+}π^{−}. In this model, the Λ_{c}^{∗+}(2593) corresponds to the lower mass Λ(1405), whereas the Λ_{c}^{∗+}(2625) corresponds to the higher mass Λ(1520) state. Although both Λ_{c}^{∗+} states can decay to Σ_{c}π, only the lower state has an allowed S-wave decay to Σ_{c}π; conservation of angular momentum and parity conservation require that the decay Λ_{c}^{∗+}(2625) → Σ_{c}π must be D-wave. For the higher mass state, since the Σ_{c}^{∗+} pole mass is kinematically inaccessible [13], the S-wave decay Σ_{c}^{∗}π is allowed only through the low-mass tail of the Σ_{c}^{∗} [14]. Therefore one expects the Λ_{c}^{∗+}(2625) to have a smaller width than the Λ_{c}^{∗+}(2593), in agreement with our measurements. The absence of any evident Σ_{c}π structure in the Λ_{c}^{∗+}(2625) band in Figure 2 is also consistent with this interpretation of the data.

It is interesting to note the splitting between the two states Λ_{c}^{∗+}(2625) and Λ_{c}^{∗+}(2593), which we measure to be 34.7±0.5±1.2 MeV/c^{2}. It is considerably less than the mass splitting between the Λ(1520) - Λ(1405), as expected both from HQET and from potential models. As anticipated in the HQET picture, it is essentially identical to the splittings observed in the mesonic sector; by comparison, M(D_{2}^{∗}(2460)^{0}) - M(D_{1}(2420)^{0}) is determined to be 35±6 MeV/c^{2} [16].

In conclusion, we have observed two peaks in the M(Λ_{c}^{+}π^{+}π^{−})-M(Λ_{c}^{+}) mass difference plot, one of which represents the first observation of a new state. The most likely explanation of these peaks is the decay of two orbitally excited Λ_{c}^{+} baryons, one decaying dominantly through an intermediate Σ_{c}π. The mass differences, M(Λ_{c}^{∗+}) - M(Λ_{c}^{+}), of these states are measured to be 307.5±0.4±1.0 and 342.2±0.2±0.5 MeV/c^{2}, respectively.

ACKNOWLEDGEMENTS

We gratefully acknowledge the effort of the CESR staff in providing us with excellent luminosity and running conditions. This work was supported by the National Science Foundation, the U.S. Dept. of Energy, the Heisenberg Foundation, the Alexander von Humboldt Stiftung, the SSC Fellowship program of TNRLC, the Natural Sciences and Engineering Research Council of Canada, and the A.P. Sloan Foundation.

REFERENCES

[1] L.Copley et al., Phys. Rev. D 20, 768 (1979).

[2] N. Isgur and G. Karl, Phys. Rev. D 18, 4187 (1978).

[3] T. DeGrand et al., Phys. Rev. D 12, 2060 (1975).

[4] A. De Rujula et al., Phys. Rev. D 12, 147 (1975).

[5] S. Capstick and N. Isgur, Phys. Rev. D 34, 2809 (1986).

[6] N. Isgur and M. Wise, Phys. Rev. Lett. 66, 1130 (1991), and references cited therein.

[7] H. Albrecht et al., (The ARGUS Collaboration), Phys. Lett. B317, 227 (1993).

[8] P. L. Frabetti et al. (The E687 Collaboration), Phys. Rev. Lett 72, 961 (1994).

[9] J. Appel, presented at the XXIII International Lepton Photon Symposium, Ithaca, NY (August, 1993), FERMILAB CONF-93-328.

[10] Y. Kubota et al., (The CLEO Collaboration), Nucl. Inst. and Methods A320, 66 (1992).

[11] P. Avery et al., (The CLEO Collaboration), Phys. Rev. Lett. 71, 2391 (1993)

[12] This number is determined by subtracting the presently tabulated value for the Λ_{c}^{+} mass (2284.6±0.9 MeV/c^{2}) from the quoted value of the Λ_{c}^{∗+} mass from reference [7].

[13] Particle Data Group, Phys. Rev. D50, 1173 (1994).

[14] P. Cho, Phys. Rev. D50, 3295 (1994).

[15] Interference effects between Σ_{c}^{++}π^{−} and the Σ_{c}^{0}π^{+} intermediate states should be small and can be neglected.

[16] J. Alexander et al., (The CLEO Collaboration), Phys. Lett. B303, 377 (1993), Y. Kubota et al. (The CLEO Collaboration), Phys. Rev. Lett. 72, 1972 (1994), and P. Avery et al. (The CLEO Collaboration), Phys. Lett. B331, 236 (1994).

[17] C. Peterson et al., Phys. Rev. D27, 105 (1983).

[18] To determine the efficiency, we generate events using the measured value for the fraction of events proceeding through a Σ_{c} intermediate state.

[19] P. Avery et al., (The CLEO Collaboration), Phys. Rev. D43, 3599 (1991).